OpenStudy (anonymous):

Two triangular roofs are similar. The ratio of the corresponding sides of these roofs is 2:3. If the altitude of the smaller roof is 4.5 feet, find the corresponding altitude of the larger roof. Round to the nearest hundredth. Help explain step by step?

Directrix (directrix):

The scale factor of two similar polygons is the same value as the ratio of any pair of corresponding linear parts of the the two polygons. By linear part, I mean a segment that can be measured with a ruler in linear units such as feet or meters.

Directrix (directrix):

In this problem, the proportion is the following: 2/3 = 4.5/x where x is the length of the altitude of the larger roof. Cross multiplying, 2x = 3 * 4.5 Solve for x and post your answer. Someone will be by to check. Okay? @Foreverisabell

OpenStudy (anonymous):

Thank you @Directrix so x = 6.75? @Directrix

Directrix (directrix):

Let me check.

Directrix (directrix):

6.75 is what I got.